what is -0.2 in exponential notation?
Answer by stanbon(75887)

what is -0.2 in exponential notation?
-----
-0.2 = -2/10 = -1/5
----
The negative means "invert".
The 5 in the denominator means "take the 5th root".
======================
Cheers,
Stan H.
========